OpenStudy (winsteria):

FAN AND MEDAL The scatterplot and trend line show a positive correlation between advertising costs and sales. Using the points (6.4, 117) and (6.6, 120), what is the slope of the trend line? 1)–15 2)–12 3)12 4)15

OpenStudy (winsteria):

@atreyu6s3x6

OpenStudy (pawanyadav):

D as 120-117/(6.6-6.4)

OpenStudy (anonymous):

okay so to find slope we use: m=y2-y1/x2-x1

OpenStudy (pawanyadav):

Which is 15

OpenStudy (winsteria):

120-117=3 6.6-6.4=0.2 3/0.2=15

OpenStudy (winsteria):

Thank you

OpenStudy (pawanyadav):

No problem

OpenStudy (anonymous):

correct good job